Vernon Manufacturing Company was started on January 1, 2018, when it acquired $84,000 cash by issuing common stock. Vernon immediately purchased office furniture and manufacturing equipment costing $7,700 and $25,300, respectively. The office furniture had an 8-year useful life and a zero salvage value. The manufacturing equipment had a $3,100 salvage value and an expected useful life of three years. The company paid $11,400 for salaries of administrative personnel and $15,100 for wages to production personnel. Finally, the company paid $15,220 for raw materials that were used to make inventory. All inventory was started and completed during the year. Vernon completed production on 4,600 units of product and sold 3,700 units at a price of $15 each in 2018. (Assume that all transactions are cash transactions and that product costs are computed in accordance with GAAP) Required a. Determine the total product cost and the average cost per unit of the inventory produced in 2018. (Round Average cost per unit to 2 decimal places.) b. Determine the amount of cost of goods sold that would appear on the 2018 income statement. (Do not round intermediate calculations.) c. Determine the amount of the ending inventory balance that would appear on the December 31, 2018, balance sheet. (Do not round intermediate calculations.) d. Determine the amount of net income that would appear on the 2018 income statement. e. Determine the amount of retained earnings that would appear on the December 31, 2018, balance sheet. f. Determine the amount of total assets that would appear on the December 31, 2018, balance sheet.
Answer #1

a. Product cost = Depreciation on manufacturing equipment + Raw material cost + wages expenses

Depreciation = ($25300 - $3100)/3 = $7400

Product cost = $7400 + $15100 + $15220 = $37720

Average cost = ($37720/4600 units) = $8.2 per unit

b. Cost of goods sold = ($8.2*3700) = $30340

c. Ending inventory = (4600 - 3700)*$8.2 = $7380

d. Invome Statement

Particulars Amount
Sales (3700*$15) $55500
Less : Cost of goods sold ($30340)
Gross Margin $25160
Less : Depreciation on office furniture ($7700/8) ($962.5)
Less : Salary and administrative expenses ($11400)
Net Income $12797.5